OpenStudy (anonymous):

Solve the system of linear equations below. 2x + 5y = 34 x + 2y = 14

OpenStudy (igreen):

(2, 6) https://www.desmos.com/calculator

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @jocelynquilesx

OpenStudy (anonymous):

Multiply bottom by two. 2x+4y=28. Subtract bottom from top. y=34-28=6 plug in to solve oth x=2

OpenStudy (anonymous):

Solve the system of linear equations below. 2x + y = 5 x + y = 4?